OpenStudy (anonymous):

A circle has a diameter of twelve millimeters. What are the circumference and area of the circle? Use three point one four for pi. Round to the nearest tenth. thirty-seven point seven millimeters; four hundred fifty-two point two square millimeters eighteen point eight millimeters; one hundred thirteen point zero square millimeters thirty-seven point seven millimeters; one hundred thirteen point zero square millimeters eighteen point eight millimeters; four hundred fifty-two point two square millimeters

OpenStudy (anonymous):

The formulas you need to know for this question are that \[A = \pi \frac{ d^{2} }{ 4 }\] and that \[C = \pi d \]

OpenStudy (anonymous):

so what do I do

OpenStudy (anonymous):

Literally, all you have to do is plug in that value for d (12 millimeters) and into the equation with A will give you the area of the circle and the equation for C will give you the circumference.

OpenStudy (anonymous):

okay thanks